# Auditing a Dutch Public Sector Risk Profiling Algorithm Using an Unsupervised Bias Detection Tool

### v2

@@ 3.1 Notation and Problem Description
Suppose that we have a total of $N$ students in the dataset, labeled $u_{1},u_{2},\ldots u_{N}$. Each row in the dataset reflects characteristics $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ (e.g., education level, age, distance to parents) of a student $u_{i}$ receiving a grant. Suppose also that we have a risk profiling algorithm $f:\mathbb{R}^{d}\rightarrow\{0,1\}$ (e.g., the risk profiling algorithm used by DUO) that takes the features of a user as input and outputs whether a user should be deemed high risk of unduly using a grant. There exists a true value of this variable of interest, denoted $y_{i}$, denoting whether the grant was, in fact, duly given (or not).
@@
We measure the degree to which an algorithm is biased by using a bias metric $M$. For a user $u_{i}$, the value of the metric $m_{i}=M(f(\boldsymbol{x}_{i}),y_{i})$ is some function of the prediction $f(\boldsymbol{x}_{i})$ and (potentially) $y_{i}$. To give some concrete examples, this could be a measure of performance, such as accuracy (requiring $f(\boldsymbol{x}_{i}),y_{i}$). Alternatively, it could measure demographic parity (only requiring $f(\boldsymbol{x}_{i})$), equality of opportunity, and so on. Without loss of generality, we assume that a larger average value for $m_{i}$ for one group compared to another group indicates bias towards the former group.
@@
For $k\subset\{1,2,\ldots,N\}$, let $U_{k}=\{u_{i}\}_{i\in k}$ be a subset of all applicants $U$. We denote by $\bar{M}(U_{k})$ the mean of the metric $m_{i}$ among all in $U_{k}$. Our goal is to test whether the difference between $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ is statistically significant. If we had access to a demographic group for which we would like to investigate bias, such as race or gender, we would simply define the subset $U_{k}$ to be different genders, race or intersections thereof. However, since the demographic group is unavailable in our setting, we leverage clustering algorithms to separate the data into subsets. Let $\mathcal{K}=\{1,...,K\}$ denote the set of clusters, and $\boldsymbol{x}_{i,k},y_{i,k},m_{i,k}$ to denote data from a user $u_{i}$ belonging to cluster $k\in\mathcal{K}$.
3.2 Clustering Algorithm
@@
In this paper, we use the Hierarchical Bias-Aware Clustering methodology from Misztal-Radecka and Indurkhya [49], hereafter referred to as “HBAC.” HBAC is an iterative clustering algorithm that produces a partition of the original dataset to isolate potentially unfairly treated groups into distinct clusters. We use this algorithm as a starting point for the DUO audit, as it represents a standard method for unsupervised bias detection in recommendation systems [21]. Additionally, we will propose methodological enhancements to refine the algorithm for real-world AI auditing applications.
@@
The HBAC algorithm first considers all observations in the dataset to belong to a single cluster and then iteratively splits the observations into distinct clusters. At each iteration, the algorithm identifies the cluster with the highest standard deviation of a specified bias metric, which serves as an indicator of imbalance between clusters. This cluster is then split into two subclusters. The process repeats for a predefined number of iterations, $max\_iterations$. Pseudocode for the algorithm is provided in Algorithm 1. Compared to the original algorithm from Misztal-Radecka and Indurkhya [49], we allow for clusters to be defined via $k$-modes in addition to $k$-means to allow clustering based on categorical and binary variables.
@@
1 Define the partition = $\{\mathcal{X}\}$
2 for $i\leftarrow 1$ to max_iterations do
3       Set $\mathcal{C}$ to be the cluster in partition with the highest standard deviation of metric $M$ among those that have not been selected in any previous iteration.
4       Split $\mathcal{C}$ into two clusters $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ using $k$-means or $k$-modes
5       if $\max(\bar{M}(\mathcal{C}^{{}^{\prime}}),\bar{M}(\mathcal{C}^{{}^{\prime\prime}}))\geq\bar{M}(\mathcal{C})\hskip 7.11317pt\land\hskip 7.11317pt|\mathcal{C}^{{}^{\prime}}|\geq n_{\mathrm{min}}\hskip 7.11317pt\land\hskip 7.11317pt|\mathcal{C}^{{}^{\prime\prime}}|\geq n_{\mathrm{min}}$  then
6             Remove $\mathcal{C}$ from partition
7             Add $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ to partition
8       end if
9
10 end for

### v3

@@ 3.1 Notation and problem description
Suppose that we have a total of $N$ students in the dataset, labeled $u_{1},u_{2},\ldots u_{N}$. Each row in the dataset reflects characteristics $\boldsymbol{x}_{i}\in\mathbb{R}^{d}$ (e.g., education level, age, distance to parents) of a student $u_{i}$ receiving a grant. Suppose also that we have a risk profiling algorithm $f:\mathbb{R}^{d}\rightarrow\{0,1\}$ (e.g., the risk profiling algorithm used by DUO) that takes the features of a user as input and outputs whether a user should be deemed high risk of unduly being given a grant. There exists a true value of this variable of interest, denoted $y_{i}$, denoting whether the grant was, in fact, duly given (or not).
@@
We measure the degree to which an algorithm is biased by using a bias metric $M$. For a user $u_{i}$, the value of the metric $m_{i}=M(f(\boldsymbol{x}_{i}),y_{i})$ is some function of the prediction $f(\boldsymbol{x}_{i})$ and (potentially) $y_{i}$. To give some concrete examples, this could be a measure of performance, such as accuracy (requiring $f(\boldsymbol{x}_{i}),y_{i}$). Alternatively, it could measure demographic parity (requiring only $f(\boldsymbol{x}_{i})$), equality of opportunity, etc. Without loss of generality, we assume that a higher average value for $m_{i}$ for one group compared to another group indicates bias towards the former group.
@@
For $k\subset\{1,2,\ldots,N\}$, let $U_{k}=\{u_{i}\}_{i\in k}$ be a subset of all applicants $U$. We denote by $\bar{M}(U_{k})$ the mean of the metric $m_{i}$ among all in $U_{k}$. Our goal is to test whether the difference between $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ is statistically significant. If we had access to a demographic group for which we would like to investigate bias, we would simply define the subset $U_{k}$ to be, for example, different genders, races, or intersections thereof. However, since the demographic group is unavailable in our setting, we leverage clustering algorithms to separate the data into subsets. Let $\mathcal{K}=\{1,...,K\}$ denote the set of clusters, and let $\boldsymbol{x}_{i,k},y_{i,k},m_{i,k}$ denote the data associated with user $u_{i}$ belonging to cluster $k\in\mathcal{K}$.
3.2 The clustering algorithm
@@
In this paper, we use the Hierarchical Bias-Aware Clustering methodology from Misztal-Radecka and Indurkhya (2021), hereafter referred to as “HBAC.” HBAC is an iterative clustering algorithm that produces a partition of the original dataset to isolate potentially unfairly treated groups into distinct clusters. We use HBAC as a starting point for the DUO audit, as it represents a standard method for unsupervised bias detection in recommendation systems (Deldjoo et al., 2024). Additionally, we propose methodological enhancements to refine the algorithm for real-world algorithmic auditing.
@@
The HBAC algorithm first considers all observations in the dataset to belong to a single cluster and then iteratively splits the observations into distinct clusters. At each iteration, the algorithm identifies the cluster with the highest standard deviation of a specified bias metric, which serves as an indicator of imbalance between clusters. This cluster is then split into two subclusters. The process repeats for a predefined number of iterations, max_iterationsmax\_iterations. The pseudocode for the algorithm is provided in Algorithm 1 in Appendix A. Compared to the original algorithm from Misztal-Radecka and Indurkhya (2021), the proposed method supports $k$-modes (enabling clustering over categorical variables) alongside $k$-means.
The original algorithm from Misztal-Radecka and Indurkhya (2021) does not specify how to assign clusters to data points that are not part of the original training dataset. We propose using the centroids of each cluster. The datapoint is assigned to the cluster whose distance from the centroid is the smallest. In the case of $k$-means and $k$-modes, the centroids are defined as, respectively, the mean and mode for each of the characteristics $\boldsymbol{x}_{i,k}$ (Huang, 1998; James et al., 2023). The Euclidean distance ($k$-means) or Hamming distance ($k$-modes) is used to determine which centroid is closer.
3.3 How to test differences in bias between clusters
In the original paper, no detailed methodology is outlined for testing whether there is a statistically significant difference in the bias metric between the clusters identified by the HBAC algorithm. It is not immediately evident that the most deviating cluster, with respect to the bias metric, is statistically significantly different from data points outside this cluster. To address this, we formulate a null hypothesis that assumes no difference in bias between the most deviating cluster and the rest of the dataset. Using a bootstrapped null-distribution, we test for differences between clusters, allowing for the construction of a valid statistical hypothesis test. We further suggest, for appropriate statistical testing, to apply sample splitting and adjustments for multiple hypothesis testing. We define the aforementioned tests in Appendix D. Following that, in Appendix E, we illustrate the test results showing that the combination of the Bonferroni correction (Hochberg and Tamhane, 1987) and sample splitting prevents us from wrongly concluding that there is a difference in the bias metric while there is none. Thus, we propose methodological improvements to the original HBAC algorithm that minimize false bias detection when using cluster analysis.
3.4 The open-source bias detection tool
The unsupervised bias detection tool is available on Github.  [^7] https://github.com/NGO-Algorithm-Audit/unsupervised-bias-detection It implements the HBAC algorithm for detecting potentially discriminated user groups, along with the methodological improvements described in Section 3.3. Currently, the library supports k-means and k-modes algorithms for splitting the data into clusters. The modular design of the tool ensures that additional splitting methods can be integrated with minimal effort. To enhance accessibility, the tool is embedded within a web application,  [^8] https://algorithmaudit.eu/technical-tools/bdt/#web-app enabling people with no programming background to use it.
4 Experiment setup: applying the bias detection tool to DUO’s risk profiling
We were granted access to ground truth labels for demographic groups in the form of aggregated statistics on students’ migration backgrounds from Statistics Netherlands (2024b). This case study is therefore well-suited to evaluate the effectiveness of the unsupervised bias detection tool in detecting bias from DUO’s risk profiling algorithm. If these ground-truth group-level data had not been available, the unsupervised bias detection tool discussed in our paper could have served as a method to examine potential bias in the DUO verification of whether a student lived at their stated address.
4.1 Description of the DUO case study
From 2012 to 2023, DUO aimed to verify that college grants were allocated to students who genuinely lived away from their parents. Some students claimed to have an independent living status to qualify for a grant, even though they did not actually reside at the address they provided (Dutch Parliament, 2024b). The process used to verify whether a student lived at the stated address is known as the CUB process (Controle Uitwonendenbeurs) which translates to “Check Non-Resident Grant.” As part of the CUB process, students were assigned a risk score ranging between 0 and 180 by a rule-based algorithm (further details are in Appendix B). This score was subsequently binned into six risk categories: very high (1), high (2), average (3), low (4), very low (5) to unknown (6). [^9] Risk category 6 includes students with a risk score of 0. Risk categories 5, 4, 3, 2, and 1 correspond to students with risk scores in the ranges of 1-19, 20-39, 40-59, 60-79, and 80-180 respectively (see Appendix B). Students in categories (1) and (2) were considered “high risk” of being unduly allocated the grant. Based on the assigned risk category, human analysts of DUO followed operational instructions to subject students to an investigation. Although students from all risk categories could potentially be investigated, the majority of investigations focused on higher-risk categories (Dutch Parliament, 2024b).
After concerns were raised regarding potential bias in the CUB process affecting students with a non-European migration background, DUO requested Statistics Netherlands to provide aggregated statistics  [^10] Through protocols established by Statistics Netherlands, the aggregated population statistics can be computed securely. Data were analyzed in a privacy-preserving manner within a secure environment. Groups are rounded to the nearest ten. Results pertaining to groups smaller than 10 people are not published. on students’ countries of origin and countries of birth  [^11] Students with a non-European migration background are defined as those who were either born outside the Netherlands and Europe (Statistics Netherlands, 2024a), or who have at least one parent born outside these regions. who received a college grant for the years 2014, 2017, 2019, 2021, and 2022 (Statistics Netherlands, 2024b). From these data, the non-European migration background of students can be retrieved in aggregate. We do not have individual-level data on students’ migration backgrounds.
This paper focuses on the years 2014 and 2019, covering a total of 298,882 students who received college grants. We specifically focus on the CUB population from 2014 ($n=248,649$) as this was the last year all types of students—MBO 1-2 and MBO 3-4 (vocational education), HBO (higher professional education), and WO (university education)—were awarded a college grant. We do not combine data from the different years since in 2019 only MBO students were eligible for a college grant. The results are replicated for the CUB 2019 dataset ($n=50,233$ students) in Appendix C.
DUO’s risk profiling algorithm differentiated on the basis of three variables. First, the type of education the student is following, categorized as MBO 1-2, MBO 3-4 (vocational education), HBO (higher professional education), and WO (university education). Second, the age of the student, categorized as: 15-18, 19-20, 21-22, 23-24, and 25-50 years. Third, the distance between the address of the parents and the student, measured through the 6-digit postal code of the registered addresses and categorized as: 0km, 1m-1km, 1-2km, 2-5km, 5-10km, 10-20km, 20-50km, 50-500km, and unknown. We note that 0km indicates that the student and their parent(s) share the same postal code but not the same house number.
In 2014, 34,050 students were assigned the risk category 6 (unknown), e.g., because the distance to their parents was unknown. We exclude these students from the clustering analysis, since the algorithm did not make a prediction on their risk of undue allocation of the grant, and these cases were deprioritized in the control process (Dutch Parliament, 2024b). Clustering is applied to the remaining dataset of $n=214,599$ students.
4.2 Applying the unsupervised bias detection tool to the CUB 2014 dataset
Data preprocessing: We transform the 3 categorical profiling characteristics to one-hot encoded variables. We define no ordinal relationship between categories. We use the $k$-modes clustering within Algorithm 1 as defined in Appendix A.
Bias metric: As stated before, our metric requires that the probability of being classified as “high risk” by the risk profiling algorithm (based on $f(x_{i})$) is equal for students with a European migration background and students with a non-European migration background. Thus, if these probabilities are equal, this condition satisfies our bias metric and corresponds to demographic parity (Dwork et al., 2012). We consider this to be an appropriate bias metric, as classifications by the risk profiling algorithm contributed to discrimination in the CUB process (Dutch Parliament, 2024a). In addition, the assessment of the adverse effect of the risk profiling algorithm on different demographic groups is a crucial aspect to investigate indirect discrimination under European non-discrimination law. An alternative definition of bias could involve examining whether college grants were duly awarded ($y_{i}$). However, ground truth labels indicating actual outcomes are only available for a small subset of students who underwent manual review by DUO (n=3,179). Consequently, we use predicted risk scores ($f(x_{i})$) instead, enabling the application to a substantially larger population (n=298,882). We emphasize that the HBAC algorithm can be used for any bias metric of interest.
Hyperparameters: The minimum number of samples per cluster is selected by iterating over 5,000, 10,000, 20,000, and 30,000 samples and selecting the one which maximizes the average Calinksi-Harabasz index (Caliński and Harabasz, 1974), [^12] This index measures the ratio of the between cluster variance over the within cluster variance. Intuitively, we want the clusters to have little variation in the bias metric within each cluster, but large variation between the clusters. This means that the data points are well separated into distinct clusters. measured via 5-fold cross-validation. We set the maximum number of iterations sufficiently high (max_iterations=1,000max\_iterations=1,000), such that Algorithm 1 in Appendix A continues until the minimum number of samples per cluster is reached. The initial centroids are determined via the method proposed by Cao et al. (2009).
Testing procedure: We split our dataset and reserve 20% of it for testing the differences in the bias metric per cluster. We fit the unsupervised bias detection tool (and determine hyperparameters) on the remaining 80% of the data. Since our bias metric is binary, we use a Pearson $\chi^{2}$-test (Pearson, 1900).
5 Bias audit results and discussion of DUO’s risk profiling algorithm
In this section, we outline the findings of our bias audit for the CUB 2014 dataset. We include results from the CUB 2019 dataset in Appendix C. Through analyzing both CUB datasets, we show that the unsupervised bias detection tool detected bias in the risk profiling algorithm that aligned with the indirect discrimination of non-European migrants found in the initial audit.
Figure 2: The percentage of students (a) deemed as “high risk” and (b) estimated as having a non-European migration background for identified clusters based on the student population of 2014, excluding students for which the risk profile was unknown ($n=214,599$). Cluster 1 is green, cluster 2 is yellow, and cluster 3 is red.
Figure 3: The percentage of students in each cluster for the three characteristics used in the risk profiling algorithm: (a) type of education, (b) age, and (c) distance to parents within the student population of 2014, excluding students for which the risk profile was unknown ($n=214,599$). For each subgroup, the average is indicated by a dashed line. Cluster 1 is green, cluster 2 is yellow, and cluster 3 is red.
5.1 Analyzing identified clusters
The unsupervised bias detection tool returns three clusters for the CUB 2014 dataset. The characteristics of these clusters are analyzed below.
Testing the difference in bias metric per cluster: We find stark differences in the bias metric across the three clusters. In the cluster with the highest bias (cluster 3) as shown in Figure 2(a), 50% of the students are classified as “high risk” by the risk profiling algorithm, versus 3% and 19% respectively in clusters 1 and 2. These differences are statistically significant at the 0.1% level, as indicated in Table 3 in Appendix D, where we report differences between $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ for the validation set separated for testing, including the $p$-values.
Figure 4: The percentage of students with a non-European migration background for the three used profiling characteristics: (a) type of education, (b) age and (c) distance to parents for students in the CUB 2014 dataset ($n=248,650$); this number is rounded to the nearest 10 and excludes students where the distance is unknown.
Student characteristics per cluster: Figure 3 illustrates the characteristics of students within each cluster. Students in cluster 3, associated with the highest average bias, are more likely to be enrolled in vocational education (MBO 1-2, MBO 3-4), be overrepresented in age groups 15-18 and 25-50, and live on average closer to their parents. In contrast, students in cluster 1, which exhibits the lowest average bias, exclusively attend university, are more likely to be 19-20 years old, and typically live far from their parents. Students in cluster 2 tend to follow higher professional education (HBO), do not display distinct patterns regarding age and live slightly more often 20-50km away from their parents.
We now examine the extent to which the identified clusters align with the non-European migration background data, using the aggregated data provided by Statistics Netherlands (2024b). [^13] The aggregated statistics are rounded to the nearest ten, hence why there are $n=248,650$ reported students instead of $248,649$. We exclude cases where the distance is unknown, since the algorithm did not apply then.
Relationship between clusters and non-European migration background: As individual-level demographic data on students’ origin are unavailable, we estimate the percentage of students with a non-European migration background per cluster using aggregate statistics. This percentage is determined using the weighted average of the proportion of students per cluster, based on the percentage of students with a non-European migration background for each combination [^14] With 4 types of education, 5 age groups and 8 distance categories, there are in total 160 possible combinations. Each possible combination is assigned to a unique cluster, e.g. if two students belong to the same combination of education level, age, and distance to parents, they belong to the same cluster. of the three characteristics (Statistics Netherlands, 2024b). There are stark differences in the proportion of students with a non-European migration background across clusters. Figure 2(b) shows that it is estimated that 41% of students in cluster 3 have a non-European migration background, compared to 17% in cluster 2 and 10% in cluster 1. These results indicate that clusters found by the unsupervised bias detection tool are related to the demographic group of interest.
Proxies for non-European migration background: To shed light on the relationship between clusters and non-European migration background, we investigate how characteristics of the clusters (and the risk profiling algorithm) serve as proxies for non-European migration background. This is illustrated in Figure 4. Students who are enrolled in vocational education (MBO 1-2, MBO 3-4), in the age groups 23-24 and 25-50, and living less than 10km away from their parents on average are more likely to have a non-European migration background, corresponding to the characteristics of students in cluster 3. In contrast, university students (WO), young students, and students living far from their parents are less likely to have a non-European migration background. These characteristics correspond to students in cluster 1; this also aligns with the risk profile logic: younger students and those living closer to their parents are assigned higher risk scores (Dutch Parliament, 2024b).
Combinations of profiling characteristics as proxies for non-European migration background: In addition to examining single characteristic relationships, we also investigate how specific combinations of characteristics act as a proxy for having a non-European migration background. Figure 5(a) highlights that young students in university education (WO) are less likely to have a non-European migration background. This result aligns with the characteristics of cluster 1, which consists exclusively of university (WO) students and contains relatively more students aged 19-20. Relatively older students (older than 23) who follow vocational education (MBO 1-2, MBO 3-4) are also more likely to have a non-European migration background.
Figure 5(b) indicates that students who follow vocational education (MBO) and live relatively close to their parents are more likely to have a non-European migration background. This pattern is reflected in cluster 1, which contains MBO 1-2 and MBO 3-4 students, who tend to live closer to their parents. Figure 5(c) shows a strong overrepresentation of students with a non-European migration background in the 15-18 age group who live 1m-1km from their parents. This pattern is somewhat reflected in cluster 3, which contains a higher proportion of 15-18 year old students and students living 1m-1km away from their parents.
5.3 Limitations
Limitations of clustering to audit algorithmic systems: The unsupervised bias detection tool alone cannot establish prohibited discrimination, as human expertise is essential to evaluate bias within the context of relevant legislation. However, the tool can serve as a starting point for such deliberative assessments. For instance, the clusters that the tool identifies could prompt further investigation into how the characteristics of individuals relate to demographic groups.
Limitations of applying the unsupervised bias detection tool to the DUO use case: There are several challenges associated with using aggregated data to evaluate the proportion of students with a non-European migration background in the clusters identified by the tool. For example, this approach requires estimating the percentage of students with a non-European migration background and prevents precise analysis of the joint relationship between the bias metric, clusters and demographic groups. However, this is an unavoidable limitation due to the sensitive nature of the data.
6 Conclusion
As demonstrated through the DUO case study, the unsupervised bias detection tool identifies clusters that correspond to groups of students found to be unequally treated by a risk profiling algorithm. The clustering results highlight proxy characteristics associated with a non-European migration background: vocational education (MBO) and living close to parents, as examined through aggregate statistics provided by Statistics Netherlands. By improving unsupervised learning methods which are behind the tool we provide, we enhance public knowledge on auditing algorithmic decision-making processes for bias, in the absence of demographic data. In addition, the unsupervised bias detection method has been implemented as an open-source tool in the form of a Python package, accompanied by a graphical interface in a web application to encourage adoption by a non-technical audience. The tool provides a scalable method that helps private entities, as well as internal and external auditors, to obtain inferences about seemingly neutral rules or practices that might embed unwanted biases. We hope that the contributions from this paper will serve the community by assisting in the detection of biases in algorithmic decision-making systems.
Appendix A The hierarchical bias-aware clustering (HBAC) algorithm
We provide pseudocode for the HBAC algorithm below and use notation as defined in Section 3.1.
@@ Appendix A The hierarchical bias-aware clustering (HBAC) algorithm
Define the partition = $\{\mathcal{X}\}$
for $i\leftarrow 1$ to max_iterations do
   Set $\mathcal{C}$ to be the cluster in partition with the highest standard deviation of metric $M$ among those that have not been selected in any previous iteration.
   Split $\mathcal{C}$ into two clusters $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ (e.g. via $k$-means or $k$-modes)
 if $\max(\bar{M}(\mathcal{C}^{{}^{\prime}}),\bar{M}(\mathcal{C}^{{}^{\prime\prime}}))\geq\bar{M}(\mathcal{C})\hskip 7.11317pt\land\hskip 7.11317pt|\mathcal{C}^{{}^{\prime}}|\geq n_{\mathrm{min}}\hskip 7.11317pt\land\hskip 7.11317pt|\mathcal{C}^{{}^{\prime\prime}}|\geq n_{\mathrm{min}}$ then
      Remove $\mathcal{C}$ from partition
      Add $\mathcal{C}^{{}^{\prime}}$ and $\mathcal{C}^{{}^{\prime\prime}}$ to partition
   end if
end for